1-) You buy a $10.000 par Treasury bill at $9,575 and sell it 60 days later for $9,675. What is your EAR?
Answer #1

Solution :- (1).

60-day return = (selling price - buying price) / Buying Price

= ( 9,675 - 9,575 ) / 9575 = 0.01045 = 1.044%

Converting it to an EAR = ( 1 + 0.01045 )(365/60) - 1

= 1.06522 - 1

= 0.0652 = 6.52%

If there is any doubt please ask in comments